OpenStudy (anonymous):

An animal dies in the forest. how long will it be until only 7/8 of the original amount of carbon 14 remains? (carbon 14 half-life = 5568 yrs.)

OpenStudy (anonymous):

@Sourav12084

OpenStudy (amistre64):

same concept as the last one we did 2^n = 7/8, logging it out we get n=log(7/8)/log(2) so n* 5568

OpenStudy (amistre64):

got a little something off there

OpenStudy (amistre64):

1/2^n = 7/8 2^n = 8/7

OpenStudy (amistre64):

n = log(8/7)/log(2) :)

OpenStudy (anonymous):

1073 yrs

OpenStudy (anonymous):

thanks!!

OpenStudy (amistre64):

all that does is change the sign in the end ... but its good to be 'conventional'

OpenStudy (amistre64):

yep, 1072.64...